Parth (parthkohli):

\[\sum_{k = 1}^{1000} 2^{(k!)!} \equiv n\pmod{1000}\]\(n\) is a three-digit number.

Parth (parthkohli):

@amistre64 Phi function?

Parth (parthkohli):

Well, \((k!)!\) would be divisible by all positive numbers under \((k!)!\)

Parth (parthkohli):

How can I proceed with that information?

OpenStudy (amistre64):

well, the sum of mods reduces each term .. might be useful. 23 = k (mod 3) k = 2; since 3(7) + 2 = 23 23 = 3+3+3+3+3+3+3+2 3+3+3+3+3+3+3+2 = k (mod 3) 0+0+0+0+0+0+0+2 = k (mod 3) k = 2

OpenStudy (amistre64):

im not sure how applicable the phi function might be, is it stated in the question by chance? or was that just an idea of yours?

Parth (parthkohli):

Yup, know that property. :-)

Parth (parthkohli):

Idea of mine.

Parth (parthkohli):

Or would it be recognizing some pattern?

OpenStudy (amistre64):

it might be patternable ... if your mod reductions simplify down to a repetition what are the first 10 terms .. or 5 terms?

OpenStudy (amistre64):

1!, 2!, 6!, 24!, 120!, 720!, .... that can be a nightmare, but might be reduceable by phis

Parth (parthkohli):

a1 = 2 = 0 (mod 1000) a2 = 4 = 4 (mod 1000) a3 = 2^(720!) = O_O (mod 2)

Parth (parthkohli):

a3 = 2^(6!) = 2^(720) sorry

Parth (parthkohli):

2^(1!) = 2 = 2 (mod 1000) 2^(2!) = 4 = 4 (mod 1000) 2^(3!) = 64 = 64 (mod 1000) 2^(4!) = 216 (mod 1000) Hmm

Parth (parthkohli):

Do you see any pattern? o_O

OpenStudy (amistre64):

not yet ... when does 2^n equal a multiple of 1000? i know 2^10 = 1024, which is 24 mod 1000

Parth (parthkohli):

2^2 = 4 4^3 = 64 64^4 has the last three digits 216 216^5 has last three digits equal to 216

Parth (parthkohli):

no, 576.

OpenStudy (amistre64):

lol, dont look at this: http://www.wolframalpha.com/input/?i=2%5E%28n%21%29%21%2C+n%3D1..3

Parth (parthkohli):

Super-exponential function! :-O

OpenStudy (amistre64):

i see a pollard factorization method that might be useful ... but might not

Parth (parthkohli):

Dag nab those fancy names.

OpenStudy (amistre64):

2^(p-1)=1 (mod p) per fermat little thrm.

Parth (parthkohli):

Yup.

OpenStudy (amistre64):

if k! = (p-1)q for some interger q, then that reduces to 1 mod p

Parth (parthkohli):

Hmm... :-|

Parth (parthkohli):

Is there anything to do now? :-\

Parth (parthkohli):

Oh.\[k! \equiv 2^{(p - 1)} \pmod{p}\]Is it so?

OpenStudy (amistre64):

thats fermats little thrm i believe

Parth (parthkohli):

Were you able to solve it then?

Parth (parthkohli):

Gotta sleep, or rather: gotta sleep on the problem. ;-) If you have any solution, feel free to post.

Parth (parthkohli):

Thanks for your time!

OpenStudy (amistre64):

me? no, i got no idea what the solution would be for this. number theory is too far behind and too little used for me to think clearly about it.

Parth (parthkohli):

Thanks anyway sire!

OpenStudy (amistre64):

ive got my elementary number theory textbook that im looking thru, and might be able to determine something in about a day :)

Parth (parthkohli):

@experimentX Any ideas on this one?

OpenStudy (experimentx):

I don't know ... but let me see.

Parth (parthkohli):

What does Mathematica compute?

Parth (parthkohli):

:-|

OpenStudy (experimentx):

overflow :(((

Parth (parthkohli):

Oh . . . what do you mean by overflow? It isn't able to compute it due to the hugeness?

OpenStudy (experimentx):

Yep ...

Parth (parthkohli):

Darn. I thought Mathematica was smart.

OpenStudy (experimentx):

Let me try finding the sum of mods first

Parth (parthkohli):

Yeah, try it one-by-one, maybe we can find some pattern. Actually we did above, but let's see what we get for the further values.

OpenStudy (experimentx):

Sum[Mod[2^((i!)!), 1000], {i, 1, 1000}]

OpenStudy (experimentx):

returns overflow in computation ... let me try on matlab.

OpenStudy (experimentx):

looks like application of Euler totient function. but let me compute it numerically first.

Parth (parthkohli):

Yup, that.

Parth (parthkohli):

But it would get harder and harder to compute \(\phi((n!)!)\) as \(n\) increases. :-|

OpenStudy (experimentx):

trust me .. I don't know that ... I am least experience with number theory.

Parth (parthkohli):

Me too :-(

OpenStudy (experimentx):

damn!! even sum of these numbers is huge digit ... returns NaN

Parth (parthkohli):

@Callisto Hi, any ideas?

Parth (parthkohli):

Try calculating modulo 1000? Just an idea...

OpenStudy (experimentx):

bug on my code .. :(((

Parth (parthkohli):

Well, it's more like pattern recognition, I guarantee you that.

OpenStudy (experimentx):

yeah i know ... my code was correct ,,, just 2^(,,(..)) is such a huge number. i guess Bigger than grahms number.

Parth (parthkohli):

But still lesser than Googol.

Parth (parthkohli):

Or is it?

OpenStudy (experimentx):

Grahams number is bigger than googol

OpenStudy (experimentx):

numerically it's a fiasco ... let me look for totient function.

Parth (parthkohli):

Fiasco indeed.

OpenStudy (experimentx):

I never used Totient function on my entire life. ... looks like worth it.

Parth (parthkohli):

Ditto!

OpenStudy (experimentx):

freaking group theory is again here.

OpenStudy (experimentx):

up until now I managed to calculate 2^1000! mod 1000 numerically.

OpenStudy (experimentx):

A very interesting property \[ a^{\phi (n)} = a^{mn} = k \mod n \] So the problem is asking, for what value of \( k \ge n \), \((k!)! = 0 \mod \phi(1000)\)

OpenStudy (experimentx):

The answer is 454 ... Amazing, Apart for Euler theorem I managed to learn Euclid algorithm, System of linear congruence, Chinese Remainder theorem, ...

Parth (parthkohli):

Is there a CRT here too?!

OpenStudy (experimentx):

oh ... just step for understanding Euler's theorem. I have never done number theory in my entire life.

Parth (parthkohli):

I see, number theory and contest mathematics combined is pretty hard!

OpenStudy (experimentx):

quite hard ... i often have difficulty doing problems on most of them.